# HyperSIGMA: Hyperspectral Intelligence Comprehension Foundation Model

### v1

@@ 5.2.1 Experiment Settings
Due to the similarities between HAD and HTD tasks, we use a similar network for HAD implementation. The primary difference is in generating coarse detection labels, where we utilize the classical HAD method RX [181]. Since HAD does not involve the target spectrum, $\mathbf{F}^{\star}$ is used for segmentation after passing through a 1$\times$1 convolution.
@@
For segmentation, HSIs are divided into 32$\times$32 patches with an overlap of 16, given that the SpatViT’s patch size is 1. The patch embedding layers’ weights are trained from scratch, with $H^{\prime}=H$ and $W^{\prime}=W$. For the SpecViT, $N_{spec}$ is set to 100. The networks are trained for 10 epochs using AdamW [165] with a learning rate of 0.00006. The batch size is set to 2 for HTD tasks and 1 for HAD tasks. The area under the ROC curve (AUC) is used as the evaluation metric.
5.2.2 Results and Analyses
TABLE VI: AUCs (%) of various methods on HTD tasks. Best and 2nd-best results are highlighted.
Dataset CEM [180] STD [182] CSCR[183] SRBBH[184] HTD-Net[185] HTD-IRN[186] CGSAL [187] SpatSIGMA HyperSIGMA
Mosaic 97.92 99.30 98.95 93.33 89.70 99.30 99.44 99.75 99.56
AVIRIS 99.83 93.98 99.35 84.43 98.80 98.98 98.98 99.78 99.83
TABLE VII: AUCs (%) of various methods on HAD tasks. Best and 2nd-best results are highlighted.
Dataset RX[181] KIFD[188] CRD [189] GTVLRR [190] Auto-AD [191] RGAE[192] GT-HAD [193] SpatSIGMA HyperSIGMA
Pavia 99.82 93.33 99.33 99.49 99.79 99.94 99.87 99.94 99.96
Cri 96.75 99.28 91.64 88.25 97.86 97.09 94.39 98.14 99.32
Figure 12: Detection probability maps predicted by various methods for HTD tasks. Top row: Mosaic [174] Bottom row: AVIRIS[175].
Figure 13: Detection probability maps predicted by various methods for HAD tasks. Top row: Pavia[176] Bottom row: Cri[177].
We compare our models, SpatSIGMA and HyperSIGMA, with existing classical and advanced HTD methods: CEM [180], STD [182], CSCR [183], SRBBH [184], HTD-Net [185], HTD-IRN [186], and CGSAL [187], as well as HAD methods including RX [181], KIFD [188], CRD [189], GTVLRR [190], Auto-AD [191], RGAE [192], and GT-HAD [193]. The results, shown in Tables VI-VII, demonstrate that our models outperform all existing methods in both HTD and HAD tasks. HyperSIGMA, with enhanced spectral information, shows further accuracy improvements over SpatSIGMA in most cases. In summary, our models achieve state-of-the-art performance in hyperspectral detection tasks, demonstrating their effectiveness of transferability.
@@ 5.2.2 Results and Analyses
We then qualitatively visualize the detection results of different methods. As shown in Figures 12-13, our models outperform others by effectively separating targets from backgrounds, accurately identifying the entire target areas, and assigning high detection confidences to these regions. However, we notice that some unrelated object regions in our results have high probabilities, especially in the HTD task. This is likely due to inaccurate pseudo labels from coarse detection technologies, which mislead the model. Additionally, our results exhibit traces of grid splicing due to sliding window-based inferences. This effect is more pronounced in HyperSIGMA than in SpatSIGMA. We hypothesize that, while fusing spectral information enhances representational ability, the resulting channel changes may increase the instability of spatial features.
Figure 14: ROC curves for various methods on HTD and HAD tasks.
To comprehensively compare the detection performance, we plot the ROC curves of different methods in Fig. 14. These results indicate that our models are competitive across various false alarm rates (FAR). In the HTD task, HyperSIGMA performs particularly well at higher FARs, effectively recognizing most target pixels. However, at lower FARs, our models struggle with irrelevant targets, consistent with the predicted map. For the HAD task, HyperSIGMA consistently outperforms other methods across nearly all FARs. In summary, these experiments demonstrate that HyperSIGMA is highly effective for hyperspectral detection tasks, highlighting the model’s versatility.
5.3 Hyperspectral Change Detection
In this section, we fine-tune pre-trained models for hyperspectral change detection, which parallels conventional remote sensing tasks using aerial RGB images. Our focus here is primarily on classical bi-temporal scenarios.
5.3.1 Experiment Settings
Dataset We tested on four common benchmark datasets: Hermiston, Farmland, Barbara, and BayArea, where are publicly available [^4] https://citius.usc.es/investigacion/datasets/hyperspectral-change-detection-dataset [^5] https://rslab.ut.ac.ir/data.
Implementation Details Following the successful practice of SST-Former [194], a transformer-based method for hyperspectral change detection, we adopt a classification network. Similarly, we select 5$\times$5 patches as input when training on Hermiston and Farmland, and 15$\times$15 patches for BayArea and Barbara. For the spatial subnetwork, the patch size of embedding layers is 1 for Hermiston and Farmland, and 2 for BayArea and Barbara, to accommodate their greater variability. The HyperSIGMA model’s architecture for this task is depicted in Fig. 15. Firstly, we merge four spectral features $\bm{V}_{i}$ from layers $i\cdot(n/4),i=1,2,3,4$ of the spectral subnetwork, obtaining $\bm{V}$, i.e., $\sum_{i=1}^{4}\bm{V}_{i}=\bm{V}$. Then, features $\bm{F}_{i}$ from layers $i\cdot(n/4),i=1,2,3,4$ of the spatial branch are separately combined with $\bm{V}$ through four SEMs. Afterward, the spatiotemporal fusion of two temporal phases is achieved by computing subtraction results, which are then passed through two convolutional layers for information integration. The final change detection result is produced using a classifier on the concatenated features.
We compared our models against several established methods: CVA [195], ISFA [196], GETNET [197], ML-EDAN [198], BIT [199], EMS-Net [200], CSA-Net [201], SST-Former [194], and GlobalMind [202]. Following the protocol of SST-Former, we randomly sampled 500 unchanged and 500 changed pixels from each dataset for training. Our models were trained for 50 epochs using a learning rate of 0.00006 and a batch size of 32. The spectral branch utilized 144 tokens to capture detailed spectral information ($N_{spec}=144$). Model performance was evaluated using OA, Kappa score, F1 score, precision, and recall. Quantitative results for the Hermiston, Farmland, BayArea, and Santa Barbara datasets are presented in Tables VIII to XI, respectively.
@@ 5.3.1 Experiment Settings
TABLE VIII: Change detection accuracy (%) of various methods on the Hermiston dataset. Best and 2nd-best results are highlighted.
@@
TABLE IX: Change detection accuracy (%) of various methods on the Farmland dataset. Best and 2nd-best results are highlighted.
@@
TABLE X: Change detection accuracy (%) of various methods on the BayArea dataset. Best and 2nd-best results are highlighted.
@@
TABLE XI: Change detection accuracy (%) of various methods on the Santa Barbara dataset. Best and 2nd-best results are highlighted.
@@
5.3.2 Results and Analyses
Our models achieve the highest OA, Kappa, and F1 scores across all datasets, demonstrating their clear superiority over other methods. Nonetheless, we observe that the precision of our models still requires improvement, as they may misclassify unchanged areas as changed. This issue might stem from using relatively small patches for change detection, as in [194]. We speculate that using larger patches to increase context perception could mitigate this problem. Fig. 16 shows the change detection results for the BayArea dataset, highlighting True Positives (white), True Negatives (black), False Positives (red), False Negatives (green), and Uncertain areas (gray). Our models detect most changes accurately with fewer false negatives compared to other methods, resulting in superior recall. Overall, our models deliver finer and more complete detection results, demonstrating HyperSIGMA’s strong feature representation for hyperspectral change detection tasks.
5.4 Hyperspectral Unmixing
After applying HyperSIGMA to high-level downstream tasks, we extended its use to low-level tasks, starting with hyperspectral unmixing. This task aims to address the complex spectral mixtures in hyperspectral data by decomposing each pixel’s spectral signature into pure spectral signatures (endmembers) and their corresponding proportions (abundances) [203, 204, 205]. This facilitates the identification and quantification of various components in each pixel.
TABLE XII: Quantitative comparison of endmember and abundance prediction performance across various methods on the Urban dataset. Best and 2nd-best results are highlighted.
@@ 5.4 Hyperspectral Unmixing
Average 0.4373 - - 0.0865 0.2952 0.1512 0.5796 0.0598 0.0584
@@
5.4.1 Experiment Settings
Dataset
We adopt the Urban dataset [^6] https://rslab.ut.ac.ir/data to evaluate the performance of HyperSIGMA for hyperspectral unmixing.
Implementation Details Following [209], we employ a simple encoder-decoder structure for hyperspectral unmixing. Concretely, assume the input and output of the network are $X\in\mathbb{R}^{H\times W\times C}$ and $\hat{X}\in\mathbb{R}^{H\times W\times C}$, we have
@@ 5.4.1 Experiment Settings
$L_{SAD}(x_{i},\hat{x}_{i})$ $=\arccos\left(\frac{x_{i}\hat{x}_{i}}{\|x_{i}\|_{2}\|\hat{x}_{i}\|_{2}}\right),$ (11)
@@
In the implementation, we use HyperSIGMA as the feature extractor, denoted as $f_{E}$. The architecture for using HyperSIGMA for hyperspectral unmixing is shown in Fig. 17. Similar to the classification task, to comprehensively understand HSI scenes, we utilize four feature maps from the spatial branch, $\bm{F}_{i}$ (where $i=1,2,3,4$), and the final layer of the spectral branch, $\bm{V}$. These maps are fused using different SEM modules and then concatenated for information aggregation. The aggregated features pass through 1$\times$1 convolutions to reduce the dimension to $C_{a}$. Finally, the abundance matrix is generated using a softmax function.
In our experiments, we extracted 64$\times$64 patches and fed them into the network, using a patch size of 2 in the spatial ViT. We set the number of tokens $N_{spec}$ to 64. The decoder used a convolutional kernel size of 1, following [213]. Based on preliminary experiments, we set $\alpha$ to 0.35 and $\beta$ to 0.1. The model was trained for 200 epochs with a batch size of 32. We used two common quantitative evaluation indices: the mean spectral angle distance (mSAD) to compare the similarity between the learned endmembers and references, and the mean squared error (MSE) to measure the quality of the obtained abundance map. Note that lower values indicate better performance.
Figure 18: Abundance prediction results of HyperSIGMA for hyperspectral unmixing on the Urban dataset.
@@
5.4.2 Results and Analyses
Eight advanced linear and nonlinear unmixing approaches were chosen for comparison. These include three classical methods: FCLS [206], ELMM [207], and SUnSAL [208], as well as recent deep learning-based methods: CNNAEU [209], CyCU-Net [210], DeepTrans [211], and EGU-Net [212]. The quantitative results in Table XII indicate that HyperSIGMA achieves the best performance for both endmembers and abundances, demonstrating its superior feature representation capability. Qualitative evaluation of our models further supports this. Fig. 18 shows that HyperSIGMA accurately characterizes the actual distributions of mixed ground objects in HSIs. Additionally, Fig. 19 illustrates that HyperSIGMA effectively captures pure endmember signatures, closely aligning with the ground truth, especially for the tree endmember. These findings highlight the potential of our models to enhance hyperspectral low-level tasks.
5.5 Hyperspectral Image Denoising
Figure 20: Diagram illustrating the application of HyperSIGMA for hyperspectral image denoising.
TABLE XIII: Quantitative comparison of different HSI denoising methods on the WDC Mall dataset. Best and 2nd-best results are highlighted.
@@ 5.5 Hyperspectral Image Denoising
Case 5 PSNR 10.984 21.933 23.452 25.558 25.966 26.741 28.097 28.061 28.103 28.503
SSIM 0.351 0.824 0.872 0.907 0.919 0.931 0.945 0.957 0.959 0.961
SAM 0.719 0.221 0.260 0.198 0.117 0.152 0.132 0.111 0.109 0.106
@@
5.5.1 Experiment Settings
Implementation Details For HSI denoising, the HyperSIGMA model’s architecture is shown in Fig. 20. First, the HSI with added noise is input into the model. In HyperSIGMA, we use only the final layer’s spatial feature and spectral feature from their respective subnetworks. These features are then fed into additional Spatial and Spectral branches, obtaining $\bm{F}$ and $\bm{V}$. Specifically, to handle the high-dimensional output features, we stack four spatial/spectral transformer layers with an embedding dimension of 191 and a patch size of 1 to reduce the dimensions. These spatial/spectral layers are trained from scratch, producing outputs that better fit the fusion stage and reconstruction decoder. The two features are then combined via an SEM module. Before reconstruction, the input image is concatenated with the spatial-spectral fusion feature. The final image restoration is achieved through a decoder with two convolutional layers.
For each image, we select a central area of size 192$\times$192 for testing, while the remaining regions are used for training. Each iteration involves selecting 64$\times$64 patches, and the patch size of the spatial subnetwork is set to 2. The comparison methods include classical model-based approaches (LLRT [220], NGMeet [221], LRTFL0 [222], and E-3DTV [223]), and recent deep learning-based methods (DS2DP [224], QRNN3D [225], and SST [226]). We use three common evaluation metrics: Peak Signal-to-Noise Ratio (PSNR), Structure SIMilarity (SSIM), and Spectral Angle Mapper (SAM). Higher PSNR and SSIM values and lower SAM values indicate better denoising performance.
@@ 5.5.2 Results and Analyses
We present the quantitative results in Table XIII and visual comparisons of spatial and spectral reconstruction performance in Fig. 21. As shown in Table XIII, our proposed methods consistently outperform other approaches. For instance, HyperSIGMA achieves a higher PSNR than the second-best method, SST [226], especially in the more challenging Cases 3, 4, and 5. Even SpatSIGMA, the simplified version using only spatial information from the pre-trained RS model, performs comparably to SST. We attribute these results to the superior feature representations by pre-training models, which can be effectively adapted to denoising through a specially designed decoder optimized for low-dimensional tasks. Fig. 21 further highlights the advantages of our models, showing minimal spatial reconstruction errors and the least deviation in spectral curves compared to Ground Truth in Case 5. In the error maps, a higher concentration of blue pixels indicates smaller errors relative to Ground Truth, while greater similarity in spectral curves signifies better reconstruction. Both SpatSIGMA and HyperSIGMA clearly outperform other methods, with HyperSIGMA performing better due to the utilization of spectral information, aligning with the quantitative results.
5.6 Hyperspectral Image Super-Resolution
@@
TABLE XIV: Quantitative comparison of different methods on the Houston dataset at 4$\times$ and 8$\times$ scale factors. Best and 2nd-best results are highlighted.
Method Scale PSNR SSIM SAM CC RMSE ERGAS
Bicubic 4$\times$ 43.027 0.961 2.545 0.974 0.009 2.909
GDRRN [227] 4$\times$ 44.296 0.973 2.535 0.976 0.007 2.470
SwinIR [228] 4$\times$ 46.097 0.981 1.946 0.986 0.006 2.004
SSPSR [229] 4$\times$ 45.602 0.978 1.965 0.985 0.006 2.138
RFSR [230] 4$\times$ 45.868 0.979 1.830 0.986 0.006 2.066
Gelin [231] 4$\times$ 45.872 0.979 1.876 0.986 0.006 2.078
MSDFormer [232] 4$\times$ 46.201 0.981 1.777 0.987 0.006 1.996
SpatSIGMA 4$\times$ 46.337 0.982 1.789 0.980 0.005 1.954
HyperSIGMA 4$\times$ 46.455 0.982 1.782 0.988 0.005 1.923
Bicubic 8$\times$ 38.108 0.899 4.671 0.918 0.015 5.123
GDRRN [227] 8$\times$ 38.259 0.909 4.905 0.914 0.014 4.914
SwinIR [228] 8$\times$ 39.401 0.920 4.059 0.937 0.013 4.333
SSPSR [229] 8$\times$ 39.284 0.916 4.267 0.935 0.013 4.421
RFSR [230] 8$\times$ 39.490 0.921 3.840 0.938 0.013 4.297
Gelin[231] 8$\times$ 39.639 0.921 3.923 0.940 0.013 4.245
MSDFormer [232] 8$\times$ 39.745 0.923 3.661 0.942 0.012 4.211
SpatSIGMA 8$\times$ 39.895 0.926 3.628 0.943 0.012 4.121
HyperSIGMA 8$\times$ 39.940 0.927 3.552 0.944 0.012 4.101
5.6.1 Experiment Settings

### v2

@@ E.1 Implementation Details
Given the similarities between HAD and HTD tasks, we use a similar network for HAD implementation (Fig. 22). The primary difference is in generating coarse detection labels, where we utilize the classical HAD method RX [83] and the highest 0.15% probabilities are labeled as pseudo detections. Since HAD does not involve the target spectrum, $\mathbf{F}^{\star}$ is used for segmentation after passing through a 1$\times$1 convolution.
@@
Experimental Hyperparameters For segmentation, HSIs are divided into 32$\times$32 patches with an overlap of 16, given that the SpatViT’s patch size is 1. The patch embedding layers’ weights are trained from scratch, with $H^{\prime}=H$ and $W^{\prime}=W$. For the SpecViT, $N_{spec}$ is set to 100. The networks are trained for 10 epochs using AdamW [218] with a learning rate of 0.00006. The batch size is set to 2 for HTD and 1 for HAD.
E.2 More Results
@@
We qualitatively visualize the detection results of different methods. As shown in Fig. 23-24, our models outperform others by effectively separating targets from backgrounds, accurately identifying the entire target areas, and assigning high detection confidences to these regions. However, we notice that some unrelated object regions in our results have high probabilities, especially in the HTD task. This is likely due to inaccurate pseudo labels from coarse detection, which mislead the model. Our results also exhibit traces of grid splicing due to sliding window-based inference. This effect is more pronounced in HyperSIGMA than in SpatSIGMA. We hypothesize that, while fusing spectral information enhances representational ability, the resulting channel changes may increase the instability of spatial features.
@@
TABLE XVIII: Change detection accuracy (%) of various methods on the Hermiston dataset. Best and 2nd-best results are highlighted.
@@
To comprehensively compare detection performance, we present the ROC curves of various methods in Fig. 25. These results show that our models remain competitive across different false alarm rates (FAR). In the HTD task, HyperSIGMA excels at higher FARs, effectively identifying most target pixels. For the HAD task, HyperSIGMA consistently outperforms other methods across nearly all FARs. Overall, these findings demonstrate HyperSIGMA’s effectiveness and versatility in hyperspectral detection tasks.
Appendix F Hyperspectral Change Detection
F.1 Implementation Details
TABLE XIX: Change detection accuracy (%) of various methods on the Farmland dataset. Best and 2nd-best results are highlighted.
@@ F.1 Implementation Details
Network Structure The HyperSIGMA model’s architecture for this task is depicted in Fig. 26. Firstly, we merge four spectral features $\bm{V}_{i}$ from layers $i\cdot(n/4),i=1,2,3,4$ of the spectral subnetwork, obtaining $\bm{V}$, i.e., $\sum_{i=1}^{4}\bm{V}_{i}=\bm{V}$. Then, features $\bm{F}_{i}$ from layers $i\cdot(n/4),i=1,2,3,4$ of the spatial branch are separately combined with $\bm{V}$ through four SEMs. Afterward, the spatiotemporal fusion of two temporal phases is achieved by computing subtraction results, which are then passed through two convolutional layers for information integration. The final change detection result is produced using a classifier on the concatenated features.
TABLE XX: Change detection accuracy (%) of various methods on the BayArea dataset. Best and 2nd-best results are highlighted.
@@
Experimental Hyperparameters We select 5$\times$5 patches as input when training on Hermiston and Farmland, and 15$\times$15 patches for BayArea and Barbara. For the spatial subnetwork, the patch size of embedding layers is 1 for Hermiston and Farmland, and 2 for BayArea and Barbara, to accommodate their greater variability.
Following the protocol of SST-Former [92], we randomly sampled 500 unchanged and 500 changed pixels from each dataset for training. To ensure a fair comparison, the training patches have a very low overlapping rate related to the testing set, aligning to [92]. Our models were trained for 50 epochs using a learning rate of 0.00006 and a batch size of 32. The spectral branch utilized 144 tokens to capture detailed spectral information ($N_{spec}=144$).
Figure 27: Change detection maps from various methods on the Hermiston dataset.
TABLE XXI: Change detection accuracy (%) of various methods on the Santa Barbara dataset. Best and 2nd-best results are highlighted.
@@
Figure 28: Change detection maps from various methods on the Farmland dataset.
Figure 30: Change detection maps from various methods on the Santa Barbara dataset.
TABLE XXII: Quantitative comparison of endmember and abundance prediction performance across various methods on the Urban dataset. Best and 2nd-best results are highlighted.
@@
Figure 31: Abundance prediction results of different methods for hyperspectral unmixing on the Urban dataset.
@@
F.2 More Results
Additional Metrics Tables XVIII-XXI present more comprehensive evaluation results beyond F1 score, encompassing OA, Kappa, precision, and recall. We observe that the precision of our models still requires improvement, as they may misclassify unchanged areas as changed. This issue might stem from using relatively small patches for change detection, as in [92]. We speculate that using larger patches to increase context perception could mitigate this problem.
Visualization Fig. 27-30 show the change detection results for the Hermiston, Farmland, Bay Area, and Santa Barbara datasets, highlighting True Positives (white), True Negatives (black), False Positives (red), False Negatives (green), and Uncertain areas (gray). Our models detect most changes accurately with fewer false negatives compared to other methods, resulting in superior recall.
Appendix G Hyperspectral Unmixing
G.1 Implementation Details
Task Formulation Assume the input and output of the network are $X\in\mathbb{R}^{H\times W\times C}$ and $\hat{X}\in\mathbb{R}^{H\times W\times C}$, we have
@@ G.1 Implementation Details
$L_{SAD}(x_{i},\hat{x}_{i})$ $=\arccos\left(\frac{x_{i}\hat{x}_{i}^{T}}{\|x_{i}\|_{2}\|\hat{x}_{i}\|_{2}}\right),$ (11)
@@
Fine-tuning Network Architecture We use HyperSIGMA as the feature extractor, denoted as $f_{E}$. The architecture for using HyperSIGMA for hyperspectral unmixing is shown in Fig. 32. Similar to the classification task, to comprehensively understand HSI scenes, we utilize four feature maps from the spatial branch, $\bm{F}_{i}$ (where $i=1,2,3,4$), and the final layer of the spectral branch, $\bm{V}$. These maps are fused using different SEM modules and then concatenated for information aggregation. The aggregated features pass through 1$\times$1 convolutions to reduce the dimension to $C_{a}$. Finally, the abundance matrix is generated using a softmax function.
Experimental settings In our experiments, we extracted 64$\times$64 patches and fed them into the network, using a patch size of 2 in the spatial ViT. We set the number of tokens $N_{spec}$ to 64. The decoder used a convolutional kernel size of 1, following [110]. Based on preliminary experiments, we set $\alpha$ to 0.35 and $\beta$ to 0.1. The model was trained for 200 epochs with a batch size of 32.
TABLE XXIII: Quantitative comparison of different HSI denoising methods on the WDC Mall dataset. Best and 2nd-best results are highlighted.
@@
Figure 34: HSI denoising results using different methods on the WDC Mall dataset for the case of in the main text. Top row: average error maps across all bands. Bottom row: spectral curves at spatial location (98, 166) of the reconstructed results.
Figure 35: Diagram illustrating the application of HyperSIGMA for HSI denoising.
G.2 More Results
Predictive Performance for Each Endmember Table XXII provides the detailed model performance for each endmember, involving abundance and endmember predictions. It can be seen that our models have advantages on almost all endmembers, especially for obvious foreground endmembers, such as Tree and Roof.
Qualitative Results Fig. 31 shows that HyperSIGMA accurately characterizes the actual distributions of mixed ground objects in HSIs. Additionally, Fig. 33 illustrates that HyperSIGMA effectively captures pure endmember signatures, closely aligning with the ground truth, especially for the tree endmember.
Appendix H Hyperspectral Image Denoising
H.1 Implementation Details
Model Architecture The HyperSIGMA model’s architecture is shown in Fig. 35. First, the HSI with added noise is input into the model. In HyperSIGMA, we use only the final layer’s spatial feature and spectral feature from their respective subnetworks. These features are then fed into additional Spatial and Spectral branches, obtaining $\bm{F}$ and $\bm{V}$. Specifically, to handle the high-dimensional output features, we stack four spatial/spectral transformer layers with an embedding dimension of 191 and a patch size of 1 to reduce the dimensions. These spatial/spectral layers are trained from scratch, producing outputs that better fit the fusion stage and reconstruction decoder. The two features are then combined via an SEM module. Before reconstruction, the input image is concatenated with the spatial-spectral fusion feature. The final image restoration is achieved through a decoder with two convolutional layers.
Experimental Details We select a central area of size 192$\times$192 for testing, while the remaining regions are used for training. Each iteration involves selecting 64$\times$64 patches, and the patch size of the spatial subnetwork is set to 2.
Different Noise Cases To fully assess the model, we also test the model under other noise cases:
@@ H.2 More Results
Visualization For the case in the main text, Fig. 34 visualizes the comparisons of spatial and spectral reconstruction performance. The results further highlight the advantages of our models, showing minimal spatial reconstruction errors and the least deviation in spectral curves compared to Ground Truth. In the error maps, a higher concentration of blue pixels indicates smaller errors relative to Ground Truth, while greater similarity in spectral curves signifies better reconstruction. Both SpatSIGMA and HyperSIGMA clearly outperform other methods, with HyperSIGMA performing better due to the utilization of spectral information, aligning with the quantitative results.
Extra Cases We present more quantitative results in Table XXIII. It can be seen that our models consistently outperform other methods, especially in the more challenging Cases 3 and 4. We attribute this to the superior feature representations by pre-training models, which can be effectively adapted to denoising through a specially designed decoder optimized for low-dimensional tasks.
Appendix I Hyperspectral Image Super-Resolution
@@
TABLE XXIV: Quantitative comparison of different methods on the Houston dataset at 4$\times$ scale factors. Best and 2nd-best results are highlighted.
Method PSNR SSIM SAM CC RMSE ERGAS
Bicubic 43.027 0.961 2.545 0.974 0.009 2.909
GDRRN [125] 44.296 0.973 2.535 0.976 0.007 2.470
SwinIR [126] 46.097 0.981 1.946 0.986 0.006 2.004
SSPSR [127] 45.602 0.978 1.965 0.985 0.006 2.138
RFSR [128] 45.868 0.979 1.830 0.986 0.006 2.066
Gelin [129] 45.872 0.979 1.876 0.986 0.006 2.078
MSDFormer [130] 46.201 0.981 1.777 0.987 0.006 1.996
SpatSIGMA 46.337 0.982 1.789 0.980 0.005 1.954
HyperSIGMA 46.455 0.982 1.782 0.988 0.005 1.923
I.1 Implementation Details